Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 02

1. Section 10.5: Odd problems in problem 17-43, 61, 65.

2. Section 10.6: Problems 29, 35, 39, 43, 49

3. Section 10.7: Odd problems in problem 37-55.

4. Find a power series that converges on (1, 3) and diverges otherwise. Do the same for (1, 3], [1, 3) and [1, 3], respectively.